OpenStudy (anonymous):

Solve the De xy''''+4y'''=0. Hint: let v=y'''

OpenStudy (anonymous):

\[-3x ^{-3}\]

OpenStudy (anonymous):

So far thats what i got after doing the subsitution. But how do i get the equations

OpenStudy (anonymous):

Letting \(v=y^{(3)}\) gives \(v'=y^{(4)}\). Subbing into the equation yields \[xv'+4v=0\] Rearranging and separating the differentials gives \[\begin{align*}xv'+4v&=0\\\\ xv'&=-4v\\\\ \frac{1}{v}\frac{dv}{dx}&=-\frac{4}{x}\\\\ \frac{dv}{v}&=-4\frac{dx}{x}\end{align*}\] Integrate and solve for \(v\), then back substitute.

OpenStudy (anonymous):

\[\ln v=-4 \ln x+k1=\ln x^{-4}+k1\]

OpenStudy (anonymous):

\[v=k2 x^{-4}=u'\]

OpenStudy (anonymous):

\[u=k2 -3x^{-3}+k3\]

OpenStudy (anonymous):

Given that \[v=k_2x^{-4}\] you would back-sub to get \[y^{(3)}=k_2x^{-4}\] and integrate until you can find an expression for \(y\).

OpenStudy (anonymous):

So then \[y^2=-3x^{-3}\] \[y=\frac{ 3 }{ 2x^2 }\]

OpenStudy (anonymous):

Is that the correct way to go about it

OpenStudy (anonymous):

Close, but you're missing the constants and a few terms that you'd get from successive integration: \[\begin{align*}y^{(3)}&=k_2x^{-4}\\ &\Downarrow\\ y''&=\int k_2x^{-4}~dx\\ &=-\frac{k_2}{3}x^{-3}+k_3\\ &\Downarrow\\ y'&=-\frac{k_2}{3}\int x^{-3}~dx+k_3\int dx\\ &=\frac{k_2}{6}x^{-2}+k_3x+k_4\\ &\Downarrow\\ y&=\int\cdots\\ &=-\frac{k_2}{12}x^{-1}+\frac{k_3}{2}x^2+k_4x+k_5 \end{align*}\]

OpenStudy (anonymous):

Okay i got this. Thanks. I got one more question. Would you help me

OpenStudy (anonymous):

\[2(x-3)^2y''+5(x-3)y'-2y=0\] Hint: Let \[t=x-3\]

OpenStudy (anonymous):

Using the sub gives \[2t^2y''+5ty'-2y=0\] This is an Euler-Cauchy equation, which assumes the solution will be proportional to \(t^r\) for some constant \(r\). \[y=t^r~~\implies~~y'=rt^{r-1}~~\implies~~y''=r(r-1)t^{r-2}\] Subbing into the equation gives \[\begin{align*}2t^2[r(r-1)t^{r-2}]+5t[rt^{r-1}]-2t^r&=0\\\\ 2r(r-1)t^r+5rt^r-2t^r&=0\\\\ 2r(r-1)+5r-2&=0 \end{align*}\] Solve for \(r\).

OpenStudy (anonymous):

Since the root are \[r=-2, \frac{ 1 }{ 2 }\] \[y(x)=c1t^{-2}+c2^{\frac{ 1 }{ 2 }}=c1(x-3)^{-2}+c2(x-3)^{\frac{ 1 }{ 2 }}\]